OpenStudy (anonymous):

How do you factor (y^2-16)/(6y^2-18y-24) I ended up with (y+4)/6(y+1) am I missing a step or did I do something wrong?

OpenStudy (asnaseer):

your answer is correct

OpenStudy (asnaseer):

\[\frac{y^2-16}{6y^2-18y-24}=\frac{(y-4)(y+4)}{6(y^2-3y-4)}=\frac{(y-4)(y+4)}{6(y-4)(y+1)}\]

OpenStudy (asnaseer):

and \((y-4)\) cancels out leaving the answer you got.

OpenStudy (anonymous):

Ok. I am using MathXL and I think I typed something wrong. Thank you!

OpenStudy (asnaseer):

yw